OpenStudy (anonymous):

Solve the equation graphically over the interval [0.2pi). Express solutions to the nearest hundredth. sin^3x-cosx+1=0

OpenStudy (anonymous):

Do you mean \[\sin ^{3} (x) - \cos (x) +1 = 0\]

OpenStudy (anonymous):

Yes, I'm new to this, and have yet to figure out how to write the equations properly.

OpenStudy (anonymous):

No problem. So this means \[\sin ^{3} (x) - \cos (x) = -1\]

OpenStudy (anonymous):

If sin^3(x) = -1 and cos(x) = 0, that would work. That happens at 270 degrees.

OpenStudy (anonymous):

It also happens when cos (x) = 1 and sin^3(x) = 0 which happens at 0 degrees.

OpenStudy (anonymous):

To find all the solutions, solve for x: \[\sin(x) = (-1 - cox(x))^{1/3}\]

OpenStudy (anonymous):

I'm a bit stumped now.